OpenStudy (anonymous):

There are (49)5 ⋅ 40 books at the library. What is the total number of books at the library? 49 445 446 1

OpenStudy (anonymous):

@Nnesha @Michele_Laino

OpenStudy (anonymous):

its raised i think thats what u call it

OpenStudy (anonymous):

4 raised to 9

OpenStudy (michele_laino):

I don't understand 49(5)*40

OpenStudy (anonymous):

Exponents?

OpenStudy (anonymous):

4 to 45

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

and 4 raised to 46

OpenStudy (anonymous):

its like this @Michele_Laino

Nnesha (nnesha):

\[\huge\rm (4^9)^5 \times 4^0\] like this ?

OpenStudy (anonymous):

YES !!!! :D

Nnesha (nnesha):

alright anything to the zero power is just equal to one \[\huge\rm (anything)^0 = 1\]

OpenStudy (anonymous):

ok yea

Nnesha (nnesha):

\[\huge\rm (x^m)^n=x^{m \times n}\] just multiply the exponents and write these dwon 'n ur notebook!

Nnesha (nnesha):

down*

OpenStudy (michele_laino):

I think that it is: \[\Large {\left( {{4^9}} \right)^5} \cdot {4^0} = {4^{45}} \cdot {4^0} = {4^{45 + 0}} = ...?\]

OpenStudy (anonymous):

ohhh so its 4 then a little 45! (idk how thats called)

Nnesha (nnesha):

type your answer choices again and use ^ for exponents

OpenStudy (michele_laino):

it is power of power

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4^9

OpenStudy (anonymous):

4^45

OpenStudy (anonymous):

4^46

OpenStudy (anonymous):

1

OpenStudy (michele_laino):

45+0=45

OpenStudy (anonymous):

thats a b c and d in order :)

OpenStudy (anonymous):

yea it is so its b??

OpenStudy (michele_laino):

yes! I think so!

OpenStudy (anonymous):

thanks guys!!

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

umm i dont know who i should give the medal to

OpenStudy (anonymous):

lol

Nnesha (nnesha):

i don't really care :=) you can give to other user trust me i won't mind

OpenStudy (anonymous):

ok:D